# A Theoretical Framework for Data Efficient Multi-Source Transfer Learning Based on Cramér-Rao Bound

### v2

@@ 4.1 Single-Source Transfer Learning
In addition, t is a scalar, ${J}({{\underline{\theta}}_{0}})$ is $d\times d$ matrix, and $\left({{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}}\right)$ is a $d-dimensional$ vector, which is the element-wise subtraction of two d-dimensional vectors ${\underline{\theta}}_{1}$ and ${\underline{\theta}}_{0}$.
@@
Compared to Theorem 4.3, Proposition 4.5 exhibits a similar mathematical form, allowing us to derive the optimal transfer quantity through a similar approach. Furthermore, we observe that as the parameter dimension $d$ increases, the K-L error measure (12) increases. This suggests that for a complex model, knowledge transfer across tasks becomes more challenging, which is consistent with the findings in (Tong et al., 2021).
@@
MSFDA(Shen et al., 2023) ResNet50 66.5 21.6 56.7 20.4 70.5 54.4 48.4 75.6 62.8 84.8 85.3 77.1
DATE(Han et al., 2023) ResNet50 - - - - - - - 75.2 60.9 85.2 84.0 76.3
M3SDA(Peng et al., 2019) ResNet101 57.2 24.2 51.6 5.2 61.6 49.6 41.5 - - - - -
@@
H-ensemble(Wu et al., 2024) ViT-S 53.4 21.3 54.4 19.0 70.4 44.0 43.8 71.8 47.5 77.6 79.1 69.0
MADA(Zhang et al., 2024) ViT-S 51.0 12.8 60.3 15.0 81.4 22.7 40.5 78.4 58.3 82.3 85.2 76.1
MADA(Zhang et al., 2024) ResNet50 66.1 23.9 60.4 31.9 75.4 52.5 51.7 72.2 64.4 82.9 81.9 75.4
MCW(Lee et al., 2019) ViT-S 54.9 21.0 53.6 20.4 70.8 42.4 43.9 68.9 48.0 77.4 86.0 70.1
WADN(Shui et al., 2021) ViT-S 68.0 29.7 59.1 16.8 74.2 55.1 50.5 60.3 39.7 66.2 68.7 58.7
@@
4.2 Multi-Source Transfer Learning
Theorem 4.6.
(proved in Appendix B.5) Given a target task ${\mathcal{T}}$ with $N_{0}$ i.i.d. samples generated from underlying model $P_{X;{\underline{\theta}}_{0}}$, and K source tasks $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$ with $N_{1},\dots,N_{K}$ i.i.d. samples generated from underlying model $P_{X;{\underline{\theta}}_{1}},\dots,P_{X;{\underline{\theta}}_{K}}$, where ${\underline{\theta}}_{0},{\underline{\theta}}_{1},\dots,{\underline{\theta}}_{K}\in\mathbb{R}^{d}$. $\hat{{\underline{\theta}}}$ is denoted as the MLE (3) based on the $N_{0}$ samples from ${\mathcal{T}}$ and $n_{1},\dots,n_{K}$ samples from $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$, where $n_{i}\in[0,N_{i}]$. Denoting $s=\sum\limits_{i=1}^{K}n_{i}$ as the total transfer quantity, and $\alpha_{i}=\frac{n_{i}}{s}$ as the proportion of different source tasks, then the K-L measure (6) can be expressed as:
In (14), $t$ is a scalar denoted as
According to Theorem 4.6, we can derive the optimal transfer quantities $n_{1}^{*},\dots,n_{K}^{*}$ by minimizing (14). Equivalently, we need to find the optimal total transfer quantity $s^{*}$ and the optimal proportion vector $\underline{\alpha}^{*}$ which minimize (14). The analytical solutions of $s^{*}$ and $\underline{\alpha}^{*}$ are difficult to acquire, and we provide a method to get their numerical solutions in Appendix D. Eventually, we can get the optimal transfer quantity of each source through $n_{i}^{*}=s^{*}\cdot\alpha_{i}^{*}$
4.3 Practical Algorithm
Along with our theoretical framework, we propose a practical algorithm, OTQMS, which is applicable to all supervised target tasks, as presented in Algorithm 1. It mainly involves two stages: (1) initializing the source task parameters $\theta_{1},...,\theta_{K}$ to compute the optimal transfer quantities and (2) jointly training the target model using a resampled dataset whose sample quantity of each source corresponds to the optimal transfer quantity derived from our theory.
Algorithm 1 OTQMS: Training
2:  Parameter: Learning rate $\eta$.
3:  Initialize: randomly initialize ${\underline{\theta}}_{0}$, $\{{\underline{\theta}}_{k}\}_{k=1}^{K}$.
4:  Output: a well-trained model $f_{{\underline{\theta}}_{0}}$ for target task.
5:  for $k=1$ to $K$ do // Initialize the source parameters
6:      repeat
7:         $L_{k}\leftarrow\frac{1}{N_{k}}\sum\limits_{i=1}^{N_{k}}\ell\left(y_{S_{k}}^{i},f_{{\underline{\theta}}_{k}}(z_{S_{k}}^{i})\right)$
8:         ${\underline{\theta}}_{k}\leftarrow{\underline{\theta}}_{k}-\eta\nabla_{{\underline{\theta}}_{k}}L_{k}$
9:      until ${\underline{\theta}}_{k}$ converges;
10:  end for $D_{train}\leftarrow D_{\mathcal{T}}=\{(z^{i},y^{i})\}_{i=1}^{N_{train}}$
11:  repeat // Joint training of the target task
12:     $L_{{\mathcal{T}}}\leftarrow\frac{1}{N_{train}}\sum\limits_{i=1}^{N_{train}}\ell\left(y^{i}_{{\mathcal{T}}},f_{{\underline{\theta}}_{0}}(z^{i})\right)$
13:     ${\underline{\theta}}_{0}\leftarrow{\underline{\theta}}_{0}-\eta\nabla_{{\underline{\theta}}_{0}}L_{{\mathcal{T}}}$
14:     $\Theta\leftarrow\left[{{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}},\dots,{{\underline{\theta}}_{K}}-{{\underline{\theta}}_{0}}\right]^{T}$
15:     ${J}({{\underline{\theta}}_{0}})\leftarrow(\nabla_{{\underline{\theta}}_{0}}L_{{\mathcal{T}}})(\nabla_{{\underline{\theta}}_{0}}L_{{\mathcal{T}}})^{T}$
16:     $(s^{*},\underline{\alpha}^{*})\leftarrow\operatorname*{arg\,min}\limits_{(s,\underline{\alpha})}\frac{d}{2}\left(\frac{1}{N_{0}+s}+\frac{s^{2}}{(N_{0}+s)^{2}}\frac{\underline{\alpha}^{T}\Theta^{T}{J}({{\underline{\theta}}_{0}})\Theta\underline{\alpha}}{d}\right)$
17:     $D_{source}\leftarrow\bigcup\limits_{k=1}^{K}\left\{D_{S_{k}}^{*}\bigg{|}D_{S_{k}}^{*}\subseteq D_{S_{k}},\;\lvert D_{S_{k}}^{*}\rvert=s^{*}\alpha_{k}^{*}\right\}$
18:     $D_{train}\leftarrow D_{source}\bigcup D_{\mathcal{T}}$
19:  until ${\underline{\theta}}_{0}$ converges;
5 Experiments
5.1 Experiments Settings
Implementation Details. We employ the ViT-Small model (Wightman, 2019), pre-trained on ImageNet-21k (Deng et al., 2009), as the backbone for all datasets. The Adam optimizer is employed with a learning rate of $1e^{-5}$. We allocate 20% of the dataset as the test set, and report the highest accuracies within 5 epoch early stops in all experiments. Following the standard few-shot learning protocol, the training data for k-shot consists of k randomly selected samples per class from the target task. All experiments are conducted on Nvidia A800 GPUs.
Baselines. For a general performance evaluation, we take SOTA works under similar settings as baselines. The scope of compared methods includes: 1) Unsupervised Methods: MSFDA (Shen et al., 2023), DATE (Han et al., 2023), M3SDA (Peng et al., 2019). 2) Few-Shot Methods Based on Model(Parameter)-Weighting: H-ensemble (Wu et al., 2024), MCW (Lee et al., 2019). 3) Few-Shot Methods Based on Sample: MADA (Zhang et al., 2024), WADN (Shui et al., 2021) 4) Source Ablating Methods: Target-Only, Single-Source-Avg/Single-Source-Best (average/best performance of single-source transfer), AllSources $\cup$ Target (all source & target data in multi-source transfer). Note that MADA (Zhang et al., 2024) leverages all unlabeled target data in conjunction with a limited amount of labeled target data, which is a hybrid approach combining unsupervised and supervised learning. Due to the page limit, we provide detailed information on the experimental design and the results of an experiment adapted to the WADN settings on the Digits dataset in Appendix C.2.
@@ 5.2 Main Result
We evaluated our algorithm, OTQMS, alongside baseline methods on the multi-source transfer learning tasks using the DomainNet and Office-Home datasets. The quantitative results are summarized in Table 2. Since the unsupervised baselines are not designed for the supervised few-shot setting, we report their original results from the respective papers for reference. We make the following observations:
@@
Data Speaks Louder Than Model Weighting. It is worth noting that on both datasets, samples-based methods utilizing both target and source samples to jointly train the model, such as WADN, MADA and OTQMS, generally outperform model(parameter)-weighting approaches which construct the target model by weighting source models, such as H-ensemble and MCW. This observation suggests that sample-based approaches offer greater advantages over model-based methods, because they can fully leverage the relevant information from the source data for the target task.
Take, But Only as Much as You Need. Comparing results in Table 2 among Target-only, AllSources $\cup$ Target, and OTQMS, we observe that OTQMS achieves the best performance in both datasets by leveraging only a subset of data selected from all available sources based on model preference. This result validates our theory and highlights the benefits of identifying the optimal balance in the bias-variance trade-off. By choosing the right quantities of samples from the source tasks, we could train the target model more accurately. Furthermore, Figure 5 shows that OTQMS also significantly reduces the training time and sample usage, which validates its superiority in terms of data efficiency.
Few-Shot Labels, Big Gains. We make a comparison of the results of unsupervised and supervised methods. While other conditions remain the same, Table 2 demonstrates that even if unsupervised methods like MSFDA and M3SDA take all the target data into account (up to 1.3×$10^{5}$ samples on Real domain of DomainNet), their performance still falls short compared to the supervised methods, which rely on only a limited number of samples (3450 samples). This illustrates the importance of having supervised information in multi-source transfer learning.
5.3 Static vs. Dynamic Transfer Quantity
Table 3: Static vs. Dynamic Transfer Quantity in OTQMS on Office-Home. Arrows will transfer from the rest tasks.
@@ 5.4 Generality across Different Shot Settings
As discussed in the theoretical analysis of Theorem 4.3, our theoretical framework is applicable to any quantity of target samples. Therefore, OTQMS exhibits shot generality, enabling it to avoid negative transfer across different shot settings. To validate this, we increase the number of shots from 5 to 100 across methods including AllSources $\cup$ Target, Target-Only, and OTQMS. As shown in Figure 4, experimental results demonstrate that OTQMS consistently outperforms other approaches across all shot settings. This highlights the generality and scalability of OTQMS in terms of data utilization.
@@ 5.6 Compatibility to Parameter-Efficient Training
To evaluate the applicability with parameter-efficient training, we used a ViT-Base model (Dosovitskiy et al., 2021) integrated with the LoRA framework (Hu et al., 2021) as the backbone. This approach significantly reduces the number of trainable parameters for downstream tasks while ensuring model quality and improving training efficiency. In the experiments, we consider the trainable low-rank matrices and the classification head as the parametric model in our theoretical framework, while treating the remaining parameters as constants. Other experimental settings are the same as default. Experiments were conducted on the Office-Home dataset. The results of Table 5 in Appendix demonstrate that our method remains effective.
5.7 Analysis on Domain-specific Transfer Quantity
To understand the domain preference of OTQMS, we visualize the proportion of each source in the selected samples for each target domain. As shown in Figure 6(b), when the target domain is Clipart, OTQMS primarily leverages samples from Real, Painting, and Sketch. In addition, Quickdraw contributes minimally to any target domain. These observations align with the findings in (Peng et al., 2019). To further clarify the selection process of OTQMS during training, we visualize it in Figure 6(a). We observe that in the Office-Home dataset, Clipart initially selects all source domains but later focuses on Art and Real World. In the DomainNet dataset, Sketch predominantly selects Clipart, Painting, and Real throughout the training process.
Refer to captionRefer to caption
(a) Domain Selection
(b) Domain preference
@@ 6 Conclusion
lmpact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Notations
      Symbol       Description       ${\mathcal{T}}$       target task       $\left\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\right\}$       source tasks       $N_{0}$        quantity of target samples       $N_{1}\cdots N_{k}$       maximum sample quantity of each source       $n_{1}\cdots n_{k}$       transfer quantity of each source       $\theta$       model parameter       $\underline{\theta}$       vectorized model parameter       $\underline{\theta}_{0}$       vectorized model parameter of target task       $\underline{\theta}_{i},i\in[1,K]$       vectorized model parameter of i-th source task,       $J(\theta)$       Fisher information (scalar) of $\theta$       $J(\underline{\theta})^{d\times d}$       Fisher information (matrix) of d-dimensional $\underline{\theta}$       $P(X;\theta)$       probability of X parameterized by $\theta$       $P_{X;\theta}$       distribution of X parameterized by $\theta$       $||\underline{x}||^{2}$       l-2 norm of vector x       $\alpha_{i}$       transfer proportion in multi-source case from i-th source task       $\underline{\alpha}$       transfer proportion vector in multi-source case whose i-th entry is $\alpha_{i}$       $s$       total transfer quantity in multi-source case       $\hat{\theta}$       estimator of $\theta$       $E_{\hat{\theta}}$       expectation of $\hat{\theta}$
@@
Appendix B Proofs
B.1 Proof of Lemma 4.2
Lemma B.1.
In the asymptotic case, the K-L measure (6) and the mean squared error have the relation as follows.
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}}\middle\|P_{X;\theta_{0}}\right)\right]$

### v3

@@ 4.1 Single-Source Transfer Learning
In addition, t is a scalar, ${J}({{\underline{\theta}}_{0}})$ is $d\times d$ matrix, and $\left({{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}}\right)$ is a ddimensionald-dimensional vector, which is the element-wise subtraction of two d-dimensional vectors ${\underline{\theta}}_{1}$ and ${\underline{\theta}}_{0}$.
@@
Compared to Theorem 4, Proposition 6 exhibits a similar mathematical form, allowing us to derive the optimal transfer quantity through a similar approach. Furthermore, we observe that as the parameter dimension $d$ increases, the K-L error measure (12) increases. This suggests that for a complex model, knowledge transfer across tasks becomes more challenging, which is consistent with the findings in [25].
4.2 Multi-Source Transfer Learning
Theorem 7.
(proved in Appendix C.5) Given a target task $\mathcal{T}$} with $N_{0}$ i.i.d. samples generated from the underlying model $P_{X;{\underline{\theta}}_{0}}$, and K source tasks $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$ with $N_{1},\dots,N_{K}$ i.i.d. samples generated from the underlying model $P_{X;{\underline{\theta}}_{1}},\dots,P_{X;{\underline{\theta}}_{K}}$, where ${\underline{\theta}}_{0},{\underline{\theta}}_{1},\dots,{\underline{\theta}}_{K}\in\mathbb{R}^{d}$. $\hat{{\underline{\theta}}}$ is denoted as the MLE (4) based on the $N_{0}$ samples from $\mathcal{T}$} and $n_{1},\dots,n_{K}$ samples from $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$, where $n_{i}\in[0,N_{i}]$. Denoting $s=\sum\limits_{i=1}^{K}n_{i}$ as the total transfer quantity, and $\alpha_{i}=\frac{n_{i}}{s}$ as the proportion of different source tasks, then the proposed measure (6) can be expressed as:
In (14), $t$ is a scalar denoted as
According to Theorem 7, we can derive the optimal transfer quantities $n_{1}^{*},\dots,n_{K}^{*}$ by minimizing (14). Equivalently, we need to find the optimal total transfer quantity $s^{*}$ and the optimal proportion vector $\underline{\alpha}^{*}$ which minimize (14). The analytical solutions of $s^{*}$ and $\underline{\alpha}^{*}$ are difficult to acquire, and we provide a method to get their numerical solutions in Appendix E. Eventually, we can get the optimal transfer quantity of each source through $n_{i}^{*}=s^{*}\cdot\alpha_{i}^{*}$ .
4.3 Practical Algorithm
Along with our theoretical framework, we propose a practical algorithm, OTQMS, which is applicable to sample-based multi-source transfer learning tasks, as presented in Algorithm 1. In OTQMS, when computing the optimal transfer quantities based on Theorem 7, we use the parameters of pretrained source models to replace $\underline{\theta}_{1},\cdots,\underline{\theta}_{K}$. Considering that the source model can be trained using sufficient labeled data, it is reasonable to use the learned parameters as a good approximation of the true underlying parameters. In contrast, the number of target data in transfer learning is often insufficient, so it is difficult to accurately estimate the true parameter $\underline{\theta}_{0}$ - the parameter of the target task model - using only the target data. Therefore, as shown in lines 5-14 of Algorithm 1, we adopt a dynamic strategy. Specifically, in the first epoch, we train a $\underline{\theta}_{0}$ using only the target data. This $\underline{\theta}_{0}$ is then used, along with Theorem 7, to determine the optimal transfer quantity from each source task, and we use random sampling to form a new resampled training dataset. Finally, we continue training $\underline{\theta}_{0}$ on this new dataset, and this procedure is repeated in each subsequent epoch to iteratively update the training dataset. This mechanism helps alleviate the inaccuracy of $\underline{\theta}_{0}$, and we also validate the effectiveness of this design in Section 5.3. In particular, we compute the matrix $J$ using the gradient of $\mathcal{L}_{train}$ in line 10, and the loss function $\ell$ is the negative log-likelihood, following a widely adopted approach in deep learning known as the empirical Fisher [16, 18].
Algorithm 1 OTQMS: Training
2:Parameter: Learning rate $\eta$.
3:Initialize: randomly initialize ${\underline{\theta}}_{0}$, use parameters of pretrained source models to initialize $\{{\underline{\theta}}_{k}\}_{k=1}^{K}$.
4:Output: a well-trained $\underline{\theta}_{0}$ for target task model $f_{{\underline{\theta}}_{0}}$ .
[…]
6:repeat // Use dynamic strategy to train the target task
7:  $\mathcal{L}_{train}\leftarrow\frac{1}{|D_{train}|}\sum\limits_{(y^{i},z^{i})\in{D_{train}}}\ell\left(y^{i},f_{{\underline{\theta}}_{0}}(z^{i})\right)$
8:  ${\underline{\theta}}_{0}\leftarrow{\underline{\theta}}_{0}-\eta\nabla_{{\underline{\theta}}_{0}}\mathcal{L}_{train}$
9:  $\Theta\leftarrow\left[{{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}},\dots,{{\underline{\theta}}_{K}}-{{\underline{\theta}}_{0}}\right]^{T}$
10:  ${J}({{\underline{\theta}}_{0}})\leftarrow(\nabla_{{\underline{\theta}}_{0}}\mathcal{L}_{train})(\nabla_{{\underline{\theta}}_{0}}\mathcal{L}_{train})^{T}$
11:  $(s^{*},\underline{\alpha}^{*})\leftarrow\operatorname*{arg\,min}\limits_{(s,\underline{\alpha})}\frac{d}{2}\left(\frac{1}{N_{0}+s}+\frac{s^{2}}{(N_{0}+s)^{2}}\frac{\underline{\alpha}^{T}\Theta^{T}{J}({{\underline{\theta}}_{0}})\Theta\underline{\alpha}}{d}\right)$
12:  $D_{source}\leftarrow\bigcup\limits_{k=1}^{K}\left\{D_{S_{k}}^{*}\bigg|D_{S_{k}}^{*}\overset{\text{rand}}{\subseteq}D_{S_{k}},\lvert D_{S_{k}}^{*}\rvert=s^{*}\alpha_{k}^{*}\right\}$
13:  $D_{train}\leftarrow D_{source}\bigcup D_{\mathcal{T}}$ // Update the trainning dataset
14:until ${\underline{\theta}}_{0}$ converges;
5 Experiments
5.1 Experiments Settings
Implementation Details. We employ the ViT-Small model [30], pre-trained on ImageNet-21k [5], as the backbone for all datasets. The Adam optimizer is employed with a learning rate of $1e^{-5}$. We allocate 20% of the dataset as the test set, and report the highest accuracies within 5 epoch early stops in all experiments. Following the standard few-shot learning protocol, the training data for k-shot consists of k randomly selected samples per class from the target task. All experiments are conducted on Nvidia A800 GPUs.
Baselines. For a general performance evaluation, we take SOTA works under similar settings as baselines. The scope of compared methods includes: 1) Unsupervised Methods: MSFDA [21], DATE [9], M3SDA [19]. 2) Few-Shot Methods Based on Model(Parameter)-Weighting: H-ensemble [31], MCW [12]. 3) Few-Shot Methods Based on Sample: MADA [32], WADN [22] 4) Source Ablating Methods: Target-Only, Single-Source-Avg/Single-Source-Best (average/best performance of single-source transfer), AllSources $\cup$ Target (all source & target data in multi-source transfer).
Note that MADA [32] leverages all unlabeled target data in conjunction with a limited amount of labeled target data, which is a hybrid approach combining unsupervised and supervised learning. Due to the page limit, we provide detailed information on the experimental design and the results of an experiment adapted to the WADN settings on the Digits dataset in Appendix D.2.
@@ 5.1 Experiments Settings
MSFDA[21] ResNet50 66.5 21.6 56.7 20.4 70.5 54.4 48.4 75.6 62.8 84.8 85.3 77.1
DATE[9] ResNet50 - - - - - - - 75.2 60.9 85.2 84.0 76.3
M3SDA[19] ResNet101 57.2 24.2 51.6 5.2 61.6 49.6 41.5 - - - - -
@@
H-ensemble[31] ViT-S 53.4 21.3 54.4 19.0 70.4 44.0 43.8 71.8 47.5 77.6 79.1 69.0
MADA[32] ViT-S 51.0 12.8 60.3 15.0 81.4 22.7 40.5 78.4 58.3 82.3 85.2 76.1
MADA[32] ResNet50 66.1 23.9 60.4 31.9 75.4 52.5 51.7 72.2 64.4 82.9 81.9 75.4
MCW[12] ViT-S 54.9 21.0 53.6 20.4 70.8 42.4 43.9 68.9 48.0 77.4 86.0 70.1
WADN[22] ViT-S 68.0 29.7 59.1 16.8 74.2 55.1 50.5 60.3 39.7 66.2 68.7 58.7
@@ 5.2 Main Result
We evaluated our algorithm, OTQMS, alongside baseline methods on the few-shot multi-source transfer learning tasks using the DomainNet and Office-Home datasets. The quantitative results are summarized in Table 2. Since the unsupervised baselines are not designed for the supervised few-shot setting, we report their original results from the respective papers for reference. We make the following observations:
@@
Data Speaks Louder Than Model Weighting. It is worth noting that on both datasets, sample-based methods utilizing both target and source samples to jointly train the model, such as WADN, MADA and OTQMS, generally outperform model(parameter)-weighting approaches which construct the target model by weighting source models, such as H-ensemble and MCW. This observation suggests that sample-based approaches offer greater advantages over model-based methods, because they can fully leverage the relevant information from the source data for the target task.
Take, But Only as Much as You Need. Comparing results in Table 2 among Target-only, AllSources $\cup$ Target, and OTQMS, we observe that OTQMS achieves the best performance in both datasets by leveraging only a subset of data selected from all available sources based on model preference. This result validates our theory. By choosing the right quantities of samples from the source tasks, we could train the target model more accurately, and we give an analysis on the domain preference of transfer quantity in Appendix F. Furthermore, Figure 4 shows that OTQMS also significantly reduces the training time and sample usage, which validates its superiority in terms of data efficiency.
[…]
Table 3: Static vs. Dynamic Transfer Quantity in OTQMS on Office-Home.
@@ 5.4 Generality across Different Shot Settings
As discussed in the theoretical analysis of Theorem 4, our theoretical framework is applicable to any quantity of target samples. Therefore, OTQMS exhibits shot generality, enabling it to avoid negative transfer across different shot settings. To validate this, we increase the number of shots from 5 to 100 across methods including AllSources $\cup$ Target, Target-Only, and OTQMS. As shown in Figure 3, experimental results demonstrate that OTQMS consistently outperforms other approaches across all shot settings. This highlights the generality and scalability of OTQMS in terms of data utilization.
@@ 5.6 Compatibility to Parameter-Efficient Training
To evaluate the applicability with parameter-efficient training, we used a ViT-Base model [6] integrated with the LoRA framework [11] as the backbone. This approach significantly reduces the number of trainable parameters for downstream tasks while ensuring model quality and improving training efficiency. In the experiments, we consider the trainable low-rank matrices and the classification head as the parametric model in our theoretical framework, while treating the remaining parameters as constants. Other experimental settings are the same as default. Experiments were conducted on the Office-Home dataset. The results of Table 6 in Appendix demonstrate that our method remains effective.
Table 4: Multi-task performance on four tasks of Office-Home.
Method Backbone Office-Home
Ar Cl Pr Rw Avg
Single-task ViT-S 66.7 62.3 87.8 68.6 71.4
OTQMS ViT-S 81.7 76.0 88.6 87.5 83.5
5.7 Compatibility to Multi-Task Learning
To demonstrate the applicability of our method to multi-task learning scenarios, we conduct experiments on the Office-Home dataset using the ViT-Small model. In multi-task learning, each task simultaneously serves as both a source and a target task. In the experiments, each task is treated as the target task in OTQMS in turn during training, while the transfer quantities from all source tasks are computed in turn. This setup enables us to evaluate how effectively OTQMS leverages information across tasks. We compare the performance of our method against single-task training to evaluate its effectiveness in Table 4.
@@ Appendix A Notations
  Symbol   Description   $\mathcal{T}$}   target task   $\left\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\right\}$   source tasks   $N_{0}$    quantity of target samples   $N_{1}\cdots N_{K}$   maximum sample quantity of each source   $n_{1}\cdots n_{K}$   transfer quantity of each source   $\theta$   model parameter   $\underline{\theta}$   vectorized model parameter   $\underline{\theta}_{0}$   vectorized model parameter of target task   $\underline{\theta}_{i},i\in[1,K]$   vectorized model parameter of i-th source task,   $J(\theta)$   Fisher information (scalar) of $\theta$   $J(\underline{\theta})^{d\times d}$   Fisher information (matrix) of d-dimensional $\underline{\theta}$   $P_{X;\theta}$   distribution of X parameterized by $\theta$   $\underline{x}||^{2}$   l-2 norm of vector x   $\alpha_{i}$   transfer proportion in multi-source case from i-th source task   $\underline{\alpha}$   transfer proportion vector in multi-source case whose i-th entry is $\alpha_{i}$   $s$   total transfer quantity in multi-source case   $\hat{\theta}$   estimator of $\theta$   $E_{\hat{\theta}}$   expectation of $\hat{\theta}$
@@
Appendix B Extended Related Work
B.1 Transfer Learning Theory
Existing theoretical works can be categorized into two groups. The first group focuses on proposing measures to quantify the similarity between the target and source tasks. Within this group, some measures have been introduced, including $l_{2}$-distance [15], optimal transport cost [3], LEEP [17], Wasserstein distance [22], and maximal correlations [12]. This work belongs to the second group focusing on developing new generalization error measures. Within this group, the measures having been introduced include $f$-divergence [10], mutual information  [2], ${\mathcal{X}}^{2}$-divergence [25], $\mathcal{H}$-score [1, 31]. However, the potential of K-L divergence as a generalization error measure has not been sufficiently explored.
B.2 Multi-source Transfer Learning
Classified by the object of transfer, existing multi-source transfer learning methods mainly focus on two types: model transfer vs sample transfer [36]. Model transfer assumes there is one or more pre-trained models on the source tasks and transfers their parameters to the target task via fine-tuning [27]. This work focuses on the latter, which is based on joint training of the source task samples with those of the target task [32, 22, 14]. Classified by the strategy of transfer, existing methods mainly focus on two types: alignment strategy and matching-based strategy [35]. Alignment strategy aims to reduce the domain shift among source and target domains [13, 34, 14]. This work is more similar to the latter, focusing on determining which source domains or samples should be selected or assigned higher weights for transfer [8, 22, 25, 31]. However, most existing works either utilize all samples from all sources or perform task-level selection, whereas this work explores a framework that optimizes the transfer quantity of each source task. Moreover, many works are restricted to specific target tasks, such as classification, which limits their task generality. In addition, many studies are mainly applicable to few-shot scenarios, and may suffer from negative transfer in non-few-shot settings, which limits their shot generality, as illustrated in Table 1.
Appendix C Proofs
C.1 Proof of Lemma 3
Lemma 8.
In the asymptotic case, the proposed measure (6) and the mean squared error have the relation as follows.
[…]